# arXiv:2603.01771v1[cs.LG] 02 Mar 2026

### v1

@@ 1 Introduction
Quantile regression. Regression tasks can require measures of uncertainty. Quantile regression (Koenker and Bassett Jr, 1978) provides a way to construct prediction intervals, but typically models target individual quantiles $\tau$, or a multi-head model outputs a fixed set of quantiles (Wen et al., 2017). This can make examining arbitrary quantiles, to tailor uncertainty bounds, computationally intensive. HTI can address this, learning the dynamics $\tau\mapsto p_{\theta_{\tau}}(y|x)$ across a desired quantile range, yielding a surrogate that can predict all intermediate quantiles ($\S$5.3).
HTI is challenging, as the dynamics $\lambda\mapsto p_{\theta_{\lambda}}(y|x)$ are typically non-linear, given complex deep learning optimisation landscapes (Ly and Gong, 2025), making simple interpolation schemes, like conditional flow matching (CFM) (Lipman et al., 2022; Liu et al., 2022; Albergo and Vanden-Eijnden, 2022), unlikely to yield feasible surrogates $(\hat{p}(y|x,\lambda))_{\lambda\in\Lambda}$. HTI requires an approach capable of capturing complex, non-Euclidean dynamics from sparse ground-truth distributions. Similar problems have been addressed in standard TI (Tong et al., 2020; Scarvelis and Solomon, 2023; Kapusniak et al., 2024; Pooladian et al., 2024), however the effects of conditions on probability paths, which is essential for HTI, are currently under-explored.
@@ 2.2 Conditional optimal transport
subject to the constraint $f(y_{0}|x)+g(y_{1}|x)\leq c(y_{0},y_{1}|x),\forall(y_{0},y_{1})\in(\mathcal{Y}_{0},\mathcal{Y}_{1})$. Enforcing this constraint with neural instantiations of $f$ and $g$ across the entire domain is challenging (Seguy et al., 2017). As such, we follow recent literature (Taghvaei and Jalali, 2019; Makkuva et al., 2020; Amos, 2022; Pooladian et al., 2024) and utilise the semi-dual formulation based on the $c$-transform (Villani and others, 2008), converting the problem into an unconstrained optimisation over a single potential $g(\cdot|x)$:
@@ 2.3 Conditional Lagrangian optimal transport
In the above, the cost function $c$ is where knowledge of system dynamics can be embedded, to shape the COT maps and paths (Asadulaev et al., 2022). The standard Euclidean cost $c(y_{0},y_{1})=\|y_{0}-y_{1}\|^{2}$, for example, deems straight paths as the most efficient. To induce more complex paths, given our assumed complex hyperparameter-dynamics, we require a cost function that is path-dependent, which motivates us to use principles from Lagrangian dynamics (Goldstein et al., 1980), bringing us to the CLOT setting. Given a smooth, time-dependent curve $q_{t}$ for $t\in[0,1]$, with time derivative $\dot{q}_{t}$, and a Lagrangian $\mathcal{L}(q_{t},\dot{q}_{t}|x)$, the action of $q$, $\mathcal{S}(q|x)$, can be determined as
@@ 3 Related works
Trajectory inference. TI (Hashimoto et al., 2016; Lavenant et al., 2021) is prominent in domains such as single-cell genomics, where destructive measurements preclude tracking individual cells over time (Macosko et al., 2015; Schiebinger et al., 2019). Successful TI relies on leveraging inductive biases to generalise beyond the sparse observed times. One typical bias is based on least-action principles—assuming that populations move between observed marginals in the most efficient way possible—naturally giving rise to OT approaches (Yang and Uhler, 2018; Schiebinger et al., 2019; Tong et al., 2020; Scarvelis and Solomon, 2023; Pooladian et al., 2024). Another potential bias invokes the manifold hypothesis (Bengio et al., 2013), which posits that data resides on a low-dimensional manifold, concentrated around the observed data (Arvanitidis et al., 2021; Chadebec and Allassonnière, 2022), to encourage inferred paths to traverse dense regions of the data space (Kapusniak et al., 2024).
Neural optimal transport. NNs have been used for OT, especially in high-dimensions where classical OT algorithms are infeasible (Makkuva et al., 2020; Korotin et al., 2021). The semi-dual OT formulation with neural parametrisations of the Kantorovich potentials and transport maps is standard (Taghvaei and Jalali, 2019; Makkuva et al., 2020; Amos, 2022; Pooladian et al., 2024). Neural COT has also been explored (Wang et al., 2024; 2025), although with fixed cost functions, and we novelly extend this to incorporate learned conditional Lagrangian costs. Our work is particularly related to Scarvelis and Solomon (2023) and Pooladian et al. (2024), which jointly learn OT cost functions and resulting transport maps from observed time marginals. We consider using more expressive forms for the cost function, involving Lagrangians with kinetic and potential energy terms, and we operate in the conditional OT setting.
Conditional generative modeling via density transport. Some generative models, such as conditional diffusion (Ho and Salimans, 2022) and CFM models (Zheng et al., 2023), operate by transporting mass from a source to a target distribution, according to some condition. They can therefore be applied to conditional TI. However, generative models focus on accurately learning the target data distribution, and they are generally unconcerned with the intermediate distributions formed along the transport paths. While some recent works utilise OT principles to achieve more efficient learning and sampling for CFM models (Tong et al., 2023; Pooladian et al., 2023), their primary objective remains high-fidelity sample generation from the target distribution.
@@ 5.2.1 Cancer therapy
To learn the $\lambda_{\text{nk}}$-induced dynamics of the policy distribution across $\lambda_{\text{nk}}\in[0,10]$, we train ground-truth policies with PPO at $\lambda_{\text{nk}}\in\{0,5,10\}$ and sample $1000$ state-action pairs from each converged policy, across a shared set of states, to act as the HTI training set. We assess the four approaches from $\S$5.1, alongside some baselines. We compare to: (1) a direct surrogate, where the target $\lambda_{\text{nk}}$, current state, and action from $\lambda_{\text{nk}}=0$ are inputs to an MLP that is trained to output actions at a given $\lambda_{\text{nk}}$ via supervised regression; (2) a CFM surrogate (Lipman et al., 2022), which learns a vector field between the distributions at $\lambda_{\text{nk}}\in\{0,5,10\}$ and generates samples by integrating actions to the desired $\lambda_{\text{nk}}$ point; (3) a metric flow matching (MFM) surrogate (Kapusniak et al., 2024) that is similar to CFM, but biases the vector field to point towards dense regions of the data space; and (4) the NLOT method of Pooladian et al. (2024). We add FiLM conditioning to these to make them appropriate for HTI.
@@ Acknowledgements
Harry Amad’s studentship is funded by Canon Medical Systems Corporation.
@@ A.1 Varying neural network robustness in dynamic noise settings
Perturbations (e.g. Gaussian noise) of magnitude $\epsilon$ added to NN training data can increase robustness during inference for noisy inputs (Goodfellow et al., 2014; Madry et al., 2017). Calibrating the training noise to that expected to be seen in deployment can lead to optimal results in terms of inference-time accuracy. The hyperparameter $\epsilon$ directly controls this trade-off: higher $\epsilon$ typically increases robustness to noisy inputs but may decrease accuracy on clean data.
@@ A.3 Varying fidelity and diversity in generative modelling
Variational Autoencoders (VAEs) (Kingma et al., 2013) are generative models that learn a latent representation of data. The $\beta$-VAE (Higgins et al., 2017) introduces a hyperparameter $\beta$ that modifies the VAE objective function by weighting the Kullback-Leibler (KL) divergence term, which acts as a regulariser on the latent space. The choice of $\beta$ critically influences the model’s behaviour:
@@ C.2.4 Model details
For the CFM surrogate model (Lipman et al., 2022), we train two flow matching models, to model the vector fields between the distributions at $\lambda_{\text{nk}}=0$ and $\lambda_{\text{nk}}=5$, and between $\lambda_{\text{nk}}=5$ and $\lambda_{\text{nk}}=10$ respectively. We base our implementation on the open source code from Lipman et al. (2024), found here: https://github.com/facebookresearch/flow_matching (CC BY-NC 4.0 License). We extend this implementation to incorporate external conditions via a FiLM layer. The hyperparameters for both of the CFM models in this surrogate model are listed in Table 11.
@@ C.3.3 Model details
$\alpha$ $0.01$ for models with $\hat{\mathcal{U}}$, 0 otherwise
@@ C.5.3 Model details
$\alpha$ $0.05$ for models with $\hat{\mathcal{U}}$, 0 otherwise

### v3

# Hyperparameter Trajectory Inference with Conditional Lagrangian Optimal Transport

@@ 1 Introduction
Quantile regression. Regression tasks can require measures of uncertainty. Quantile regression (Koenker and Bassett Jr, 1978) provides a way to construct prediction intervals, but typically models target individual quantiles $\tau$, or a multi-head model outputs a fixed set of quantiles (Wen et al., 2018). This can make examining arbitrary quantiles, to tailor uncertainty bounds, computationally intensive. HTI can address this, learning the dynamics $\tau\mapsto p_{\theta_{\tau}}(y|x)$ across a desired quantile range, yielding a surrogate that can predict all intermediate quantiles ($\S$5.3).
HTI is challenging, as the dynamics $\lambda\mapsto p_{\theta_{\lambda}}(y|x)$ are typically non-linear, given complex deep learning optimisation landscapes (Ly and Gong, 2025), making simple interpolation schemes, like conditional flow matching (CFM) (Lipman et al., 2023; Liu et al., 2023; Albergo and Vanden-Eijnden, 2023), unlikely to yield feasible surrogates $(\hat{p}(y|x,\lambda))_{\lambda\in\Lambda}$. HTI requires an approach capable of capturing complex, non-Euclidean dynamics from sparse ground-truth distributions. Similar problems have been addressed in standard TI (Tong et al., 2020; Scarvelis and Solomon, 2023; Kapusniak et al., 2024; Pooladian et al., 2024), however the effects of conditions on probability paths, which is essential for HTI, are currently under-explored.
@@ 2.2 Conditional optimal transport
subject to the constraint $f(y_{0}|x)+g(y_{1}|x)\leq c(y_{0},y_{1}|x),\forall(y_{0},y_{1})\in(\mathcal{Y}_{0},\mathcal{Y}_{1})$. Enforcing this constraint with neural instantiations of $f$ and $g$ across the entire domain is challenging (Seguy et al., 2018). As such, we follow recent literature (Makkuva et al., 2020; Amos, 2023; Pooladian et al., 2024) and utilise the semi-dual formulation based on the $c$-transform (Villani and others, 2008), converting the problem into an unconstrained optimisation over a single potential $g(\cdot|x)$:
@@ 2.3 Conditional Lagrangian optimal transport
In the above, the cost function $c$ is where knowledge of system dynamics can be embedded, to shape the COT maps and paths (Asadulaev et al., 2024). The standard Euclidean cost $c(y_{0},y_{1})=\|y_{0}-y_{1}\|^{2}$, for example, deems straight paths as the most efficient. To induce more complex paths, given our assumed complex hyperparameter-dynamics, we require a cost function that is path-dependent, which motivates us to use principles from Lagrangian dynamics (Goldstein et al., 1980), bringing us to the CLOT setting. Given a smooth, time-dependent curve $q_{t}$ for $t\in[0,1]$, with time derivative $\dot{q}_{t}$, and a Lagrangian $\mathcal{L}(q_{t},\dot{q}_{t}|x)$, the action of $q$, $\mathcal{S}(q|x)$, can be determined as
@@ 3 Related works
Trajectory inference. TI (Hashimoto et al., 2016; Lavenant et al., 2021) is prominent in domains such as single-cell genomics, where destructive measurements preclude tracking individual cells over time (Macosko et al., 2015; Schiebinger et al., 2019). Successful TI relies on leveraging inductive biases to generalise beyond the sparse observed times. One typical bias is based on least-action principles—assuming that populations move between observed marginals in the most efficient way possible—naturally giving rise to OT approaches (Yang and Uhler, 2019; Schiebinger et al., 2019; Tong et al., 2020; Scarvelis and Solomon, 2023; Pooladian et al., 2024). Another potential bias invokes the manifold hypothesis (Bengio et al., 2013), which posits that data resides on a low-dimensional manifold, concentrated around the observed data (Arvanitidis et al., 2022; Chadebec and Allassonnière, 2022), to encourage inferred paths to traverse dense regions of the data space (Kapusniak et al., 2024).
Neural optimal transport. NNs have been used for OT, especially in high-dimensions where classical OT algorithms are infeasible (Makkuva et al., 2020; Korotin et al., 2021). The semi-dual OT formulation with neural parametrisations of the Kantorovich potentials and transport maps is standard (Makkuva et al., 2020; Amos, 2023; Pooladian et al., 2024). Neural COT has also been explored (Wang et al., 2024; 2025), although with fixed cost functions, and we novelly extend this to incorporate learned conditional Lagrangian costs. Our work is particularly related to Scarvelis and Solomon (2023) and Pooladian et al. (2024), which jointly learn OT cost functions and resulting transport maps from observed time marginals. We consider using more expressive forms for the cost function, involving Lagrangians with kinetic and potential energy terms, and we operate in the conditional OT setting.
Conditional generative modeling via density transport. Some generative models, such as conditional diffusion (Ho and Salimans, 2022) and CFM models (Zheng et al., 2023), operate by transporting mass from a source to a target distribution, according to some condition. They can therefore be applied to conditional TI. However, generative models focus on accurately learning the target data distribution, and they are generally unconcerned with the intermediate distributions formed along the transport paths. While some recent works utilise OT principles to achieve more efficient learning and sampling for CFM models (Tong et al., 2024; Pooladian et al., 2023), their primary objective remains high-fidelity sample generation from the target distribution.
@@ 5.2.1 Cancer therapy
To learn the $\lambda_{\text{nk}}$-induced dynamics of the policy distribution across $\lambda_{\text{nk}}\in[0,10]$, we train ground-truth policies with PPO at $\lambda_{\text{nk}}\in\{0,5,10\}$ and sample $1000$ state-action pairs from each converged policy, across a shared set of states, to act as the HTI training set. We assess the four approaches from $\S$5.1, alongside some baselines. We compare to: (1) a direct surrogate, where the target $\lambda_{\text{nk}}$, current state, and action from $\lambda_{\text{nk}}=0$ are inputs to an MLP that is trained to output actions at a given $\lambda_{\text{nk}}$ via supervised regression; (2) a CFM surrogate (Lipman et al., 2023), which learns a vector field between the distributions at $\lambda_{\text{nk}}\in\{0,5,10\}$ and generates samples by integrating actions to the desired $\lambda_{\text{nk}}$ point; (3) a metric flow matching (MFM) surrogate (Kapusniak et al., 2024) that is similar to CFM, but biases the vector field to point towards dense regions of the data space; and (4) the NLOT method of Pooladian et al. (2024). We add FiLM conditioning to these to make them appropriate for HTI.
@@ Acknowledgements
Harry Amad’s studentship is funded by Canon Inc.. This work was supported by Azure sponsorship credits granted by Microsoft’s AI for Good Research Lab.
@@ A.1 Varying neural network robustness in dynamic noise settings
Perturbations (e.g. Gaussian noise) of magnitude $\epsilon$ added to NN training data can increase robustness during inference for noisy inputs (Goodfellow et al., 2015; Madry et al., 2018). Calibrating the training noise to that expected to be seen in deployment can lead to optimal results in terms of inference-time accuracy. The hyperparameter $\epsilon$ directly controls this trade-off: higher $\epsilon$ typically increases robustness to noisy inputs but may decrease accuracy on clean data.
@@ A.3 Varying fidelity and diversity in generative modelling
Variational Autoencoders (VAEs) (Kingma and Welling, 2014) are generative models that learn a latent representation of data. The $\beta$-VAE (Higgins et al., 2017) introduces a hyperparameter $\beta$ that modifies the VAE objective function by weighting the Kullback-Leibler (KL) divergence term, which acts as a regulariser on the latent space. The choice of $\beta$ critically influences the model’s behaviour:
@@ C.2.4 Model details
For the CFM surrogate model (Lipman et al., 2023), we train two flow matching models, to model the vector fields between the distributions at $\lambda_{\text{nk}}=0$ and $\lambda_{\text{nk}}=5$, and between $\lambda_{\text{nk}}=5$ and $\lambda_{\text{nk}}=10$ respectively. We base our implementation on the open source code from Lipman et al. (2024), found here: https://github.com/facebookresearch/flow_matching (CC BY-NC 4.0 License). We extend this implementation to incorporate external conditions via a FiLM layer. The hyperparameters for both of the CFM models in this surrogate model are listed in Table 11.
@@ C.3.3 Model details
$\alpha$ $0.001$ for models with $\hat{\mathcal{U}}$, 0 otherwise
@@ C.5.3 Model details
$\alpha$ $0.01$ for models with $\hat{\mathcal{U}}$, 0 otherwise